# On the Approach Towards Equilibrium Through Momentum-Dependent Relaxation: Insights from Evolution of the Moments in Kinetic Theory

### v2

@@ II.2.1 Basic definitions and properties
Multiplying Eq.(10) with $E_{p}^{n}P_{2l}({\text{cos}}\theta)$ and integrating over $dP$, we finally obtain the equation of motion governing the moment evolutions,
$\frac{\partial}{\partial\tau}\rho_{n,l}\leavevmode\nobreak\ +\leavevmode\nobreak$ $\frac{1}{\tau}\left[{\cal{P}}(n,l)\rho_{n,l-1}+{\cal{Q}}(n,l)\rho_{n,l}+{\cal{R}}(n,l)\rho_{n,l+1}\right]$
@@
with,
${\cal{P}}(n,l)=$ $2l\frac{(2l-1)}{(4l-1)}\frac{(n+2l)}{(4l+1)}\leavevmode\nobreak\ ,$
${\cal{Q}}(n,l)=$ $\frac{2}{3}+\frac{n(8l^{2}+4l-1)}{(4l-1)(4l+3)}+\frac{2l(2l+1)}{3(4l-1)(4l+3)}\leavevmode\nobreak\ ,$
${\cal{R}}(n,l)=$ $(n-2l-1)\frac{(2l+1)}{(4l+1)}\frac{(2l+2)}{(4l+3)}\leavevmode\nobreak\ .$ (12)
$(4l+1)xP_{2l}(x)=(2l+1)P_{2l+1}(x)+2l\leavevmode\nobreak\ P_{2l-1}(x)\leavevmode\nobreak\ ,$ (13)
$\partial P_{2l}(x)/\partial x=2l\left\{P_{2l-1}(x)-xP_{2l}(x)\right\}/(1-x^{2})\leavevmode\nobreak\ .$ (14)
$\frac{\partial}{\partial\tau}\rho_{n,l}\leavevmode\nobreak\ +\leavevmode\nobreak$ $\frac{1}{\tau}\left[{\cal{P}}\rho_{n,l-1}+{\cal{Q}}\rho_{n,l}+{\cal{R}}\rho_{n,l+1}\right]$
$=-\frac{1}{\tau_{R}}\left[\rho_{n,l}-\rho_{n,l}^{eq}\right]\leavevmode\nobreak\ ,$ (15)
II.2.2 Collision kernel with the momentum dependent relaxation time approximation
Contrary to the existing studies, in the current analysis, we are considering the relaxation time of single particle distribution as a function of the particle momenta (explicitly single particle energy $E_{p}$) as $\tau_{R}=\tau_{R}^{E}(E_{p})$. Accordingly, while treating the collision term in the moment equation (11), we use the novel relaxation time collision operator [33] that conserves both particle number and energy-momentum irrespective of the momentum-dependence considered in $\tau^{E}_{R}(E_{p})$ and the hydrodynamic frame chosen,
${\cal{L}}[\phi_{p}]=$ $-\frac{E_{p}}{\tau^{E}_{R}}f_{p}^{eq}\Bigg{[}\phi_{p}-P_{0}\frac{{\langle\frac{E_{p}}{\tau_{R}^{E}}P_{0}\phi_{p}\rangle}_{0}}{{\langle\frac{E_{p}}{\tau_{R}^{E}}P_{0}P_{0}\rangle}_{0}}$
$-P_{1}\frac{{\langle\frac{E_{p}}{\tau_{R}^{E}}P_{1}\phi_{p}\rangle}_{0}}{{\langle\frac{E_{p}}{\tau_{R}^{E}}P_{1}P_{1}\rangle}_{0}}-p^{\langle\mu\rangle}\frac{{\langle\frac{E_{p}}{\tau_{R}^{E}}p_{\langle\mu\rangle}\phi_{p}\rangle}_{0}}{\frac{1}{3}{\langle\frac{E_{p}}{\tau_{R}^{E}}p_{\langle\nu\rangle}p^{\langle\nu\rangle}\rangle}_{0}}\Bigg{]}\leavevmode\nobreak\ ,$ (16)
@@ II.2.2 Collision kernel with the momentum dependent relaxation time approximation
With the given prescription, we denote the right hand side of Eq.(11) (moment of the linearized collision term) as $C_{n,l,\Lambda}$ such that the moment evolution equation becomes,
$\leavevmode\nobreak\ \leavevmode\nobreak\ +\leavevmode\nobreak\ \leavevmode\nobreak\ C_{n,l,\Lambda}\leavevmode\nobreak\ .$
To proceed further, we decompose the momentum-dependent relaxation time $\tau^{E}_{R}(E_{p})$ into a momentum-independent part $\tau_{R}$ (precisely a function of temperature $T$) and a part purely a function of single particle energy $E_{p}$ with the exponent $\Lambda$ being a positive number [34],
$\tau^{E}_{R}(E_{p})=\tau_{R}(T)E_{p}^{\Lambda}\leavevmode\nobreak\ .$ (18)
$\tau_{R}(T)={\frac{\eta}{s}}\frac{5!}{(4+\Lambda)!}\frac{1}{T^{1+\Lambda}}\leavevmode\nobreak\ .$ (19)
From now on, whenever the notation $\tau_{R}$ has been mentioned either in the text or in the figures of the result section, it indicates this thermal part and must not be confused with the total relaxation time $\tau_{R}^{E}$.
Using expression (18), the moment of the collision term $C_{n,l,\Lambda}$ becomes,
$C_{n,l,\Lambda}=$ $-\frac{1}{\tau_{R}}\Bigg{[}\left\{\rho_{n-\Lambda,l}-\delta_{l0}\leavevmode\nobreak\ \rho^{eq}_{n-\Lambda,l}\right\}-A\leavevmode\nobreak\ \delta_{l0}\leavevmode\nobreak\ \rho^{eq}_{n-\Lambda,0}$
$-B\leavevmode\nobreak\ \delta_{l0}\leavevmode\nobreak\ \left\{\rho^{eq}_{n-\Lambda,0}-\frac{\rho^{eq}_{1-\Lambda,0}}{\rho^{eq}_{2-\Lambda,0}}\rho^{eq}_{n-\Lambda+1,0}\right\}\Bigg{]}\leavevmode\nobreak\ ,$ (20)
with,
$A=$ $\frac{\rho_{1-\Lambda,0}-\rho^{eq}_{1-\Lambda,0}}{\rho^{eq}_{1-\Lambda,0}}\leavevmode\nobreak\ ,$ (21)
$B=$ $\frac{\rho_{1-\Lambda,0}-\frac{\rho^{eq}_{1-\Lambda,0}}{\rho^{eq}_{2-\Lambda,0}}\rho_{2-\Lambda,0}}{\rho^{eq}_{3-\Lambda,0}\left(\frac{\rho^{eq}_{1-\Lambda,0}}{\rho^{eq}_{2-\Lambda,0}}\right)^{2}-\rho^{eq}_{1-\Lambda,0}}\leavevmode\nobreak\ .$ (22)
With the values of with $l=0$ and $n=1,2$ it can observed that $C_{n,l,\Lambda}$ readily gives zero preserving the collisional invariant property. The resulting equations give the well known conservation equations for particle number $(n)$ and energy density $(\epsilon)$ for a system as follows,
$\frac{\partial n}{\partial\tau}+\frac{n}{\tau}=0\leavevmode\nobreak\ ,$ (23)
$\frac{\partial\epsilon}{\partial\tau}+\frac{4}{3}\frac{\epsilon}{\tau}=-\frac{2}{3}\frac{P_{L}-P_{T}}{\tau}\leavevmode\nobreak\ ,$ (24)
where we identify the moments as, $n=\rho_{1,0}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \epsilon=\rho_{2,0}$ and $\rho_{2,1}=P_{L}-P_{T}=$ pressure anisotropy.
After some tedious algebra, the collision moment $C_{n,l,\Lambda}$ can be written in the following simplified form,
$C_{n,l,\Lambda}=$ $-\frac{1}{\tau_{R}}\Big{[}\rho_{n-\Lambda,l}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ +T^{n-2}K(n,2,\Lambda)C(n,\Lambda)\rho_{2-\Lambda,l}\Big{\}}\Big{]}\leavevmode\nobreak\ ,$ (25)
with the following two functions introduced,
$K(n,m,\Lambda)=\frac{\Gamma(n-\Lambda+2)}{\Gamma(m-\Lambda+2)}\leavevmode\nobreak\ ,$ (26)
$C(n,\Lambda)=\frac{1-K(1,n,\Lambda)K(n+1,2,\Lambda)}{1-K(1,2,\Lambda)K(3,2,\Lambda)}\leavevmode\nobreak\ ,$ (27)
such that $C(1,\Lambda)=0$, $C(2,\Lambda)=1$ and $K(n,n,\Lambda)=1$. This makes energy and number conservation explicit as $C_{n,l,\Lambda}$ is zero for $n=1,2$ and $l=0$. When $\Lambda$ is zero, the moments are coupled to each other only through the $l$ indices while for any non-zero $\Lambda$, each $(n,l)$ moment is coupled to the corresponding $(n-\Lambda,l)$ moment via the collision term $C_{n,l,\Lambda}$ (see Eq.(25)). This creates an infinite ladder of coupling towards the lower $n$ moments.
For $\text{For\leavevmode\nobreak\ \leavevmode\nobreak\ }\leavevmode\nobreak\ l\neq 0\leavevmode\nobreak\ ,$
$\frac{\partial}{\partial\tau}\rho_{n,l}+\frac{1}{\tau}\left[{\cal{P}}\rho_{n,l-1}+{\cal{Q}}\rho_{n,l}+{\cal{R}}\rho_{n,l+1}\right]=-\frac{1}{\tau_{R}}\rho_{n-\Lambda,l}\leavevmode\nobreak\ ,$ (28)
For $\text{For\leavevmode\nobreak\ \leavevmode\nobreak\ }\leavevmode\nobreak\ l=0\leavevmode\nobreak\ ,$
$\leavevmode\nobreak\ \leavevmode\nobreak\ -\frac{1}{\tau_{R}}\Big{[}\rho_{n-\Lambda,0}-T^{n-1}K(n,1,\Lambda)\left\{1-C(n,\Lambda)\right\}\rho_{1-\Lambda,0}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -T^{n-2}K(n,2,\Lambda)C(n,\Lambda)\rho_{2-\Lambda,0}\Big{]}\leavevmode\nobreak\ .$ (29)
Eq.(28) is structurally not very different from the momentum-independent case (15) apart from the $\Lambda$ factor inclusion in the $n$ index of the moment to the source term at the right hand side (however, that change the dynamics anyway). But it is the $l=0$ moment in (29) that bears the effect of MDRTA the most via the $C(n,\Lambda)$ and $K(m,n,\Lambda)$ functions. It again carries its effect recursively through the couplings of moment hierarchy as mentioned earlier.
II.2.3 Initial Conditions
We consider a conformal system with particle mass $m=0$ at an initial temperature $T(\tau_{0})=1$ GeV and initial chemical potential $\mu(\tau_{0})=0$. We initialize the moments $\rho_{n,l}(\tau_{0})$ to their equilibrium values $\rho^{eq}_{n,l}$. Unlike momentum-independent relaxation time, systems with larger viscosity cannot be probed by considering smaller values of initial rescaled time. In this study, we choose two values of the initial viscosity to entropy ratio $\eta_{0}/s_{0}=0.2,0.02$. To make a comparison with the previous works [27], we then choose the initial time $\tau_{0}$ such that the scaled time is $\tau_{0}/\tau_{R}(\tau_{0})=1.0,0.1$ so that the initial parameter set matches the case for $\Lambda=0$. We then solve the coupled equations in (17) by truncating the moment equations at $n-5\Lambda$ and $l_{\text{max}}=50$ using the RK4 algorithm.
II.2.4 Computational Details
The moment evolution equation (17) is a set of infinitely coupled non-linear differential equations. To compute the moments, for practical purposes we need to truncate these equations at some finite order. The proper truncation procedure for the momentum-dependent relaxation system is non-trivial due to the coupling of moments in both $n$ and $l$ indices. In a previous work by de Brito et.al, [27] the $l$ moments were truncated by setting $\rho_{n,l}=0$ for some $l\geq l_{max}$, note that due to momentum independent relaxation time there was no $n$ coupling. This assumes that the contribution of the $l$ moments beyond $l_{max}$ can be ignored compared to the lower $l$ moments. But this type of truncation scheme no longer holds true for later times, when the $l$ moments decay and become comparable in value with respect to each other. The situation is elaborated below by analyzing different terms of Eq.(17). The free streaming evolution equation is given by,
$\frac{\partial}{\partial\tau}\rho_{n,l}=$ $-\frac{1}{\tau}{\cal{Q}}(n,l)\leavevmode\nobreak\ \rho_{n,l}$
$-\frac{1}{\tau}\left[{\cal{P}}(n,l)\leavevmode\nobreak\ \rho_{n,l-1}+{\cal{R}}(n,l)\leavevmode\nobreak\ \rho_{n,l+1}\right]\,.$ (30)
In the first term on the right hand side of Eq.(30), ${\cal{Q}}>0$, indicating that this is a decay term. The terms ${\cal{P}}$ and ${\cal{R}}$ have opposite signs while $\rho_{n,l-1}$ and $\rho_{n,l+1}$ share the same sign. On the other hand $\rho_{n,l}$ alternates sign with $l$, meaning that $\rho_{n,l}$ has opposite sign of $\rho_{n,l-1}$ and $\rho_{n,l+1}$. Initially, all the moments except $l=0$ are zero. It is observed that for the given initial condition, $\rho_{n,l+1}\ll\rho_{n,l-1}$ during the early time and the ${\cal{P}}(n,l)\leavevmode\nobreak\ \rho_{n,l-1}$ term is non-zero and contributes to a coupling between $l$ modes. At late time as the system approaches towards equilibrium, we have $\rho_{n,l+1}\sim\rho_{n,l-1}$, and eventually the last two terms on the right-hand side of Eq.(30) cancel each other. Therefore, for a large enough $\tau$, the moment $\rho_{n,l}$ decays to zero.
Now, let us consider the case with a non-zero collision kernel. The truncation for the momentum-dependent case is even more complex as there is coupling in both $l$ and $n$ indices (see Eq.(25)). A non-zero $\Lambda$ fundamentally changes the type of the coupled differential equation. To understand this, let us rewrite the set of coupled differential equations in the following matrix representation
$\frac{d}{d\tau}\vec{\rho}=-\frac{1}{\tau}\mathbf{F}\vec{\rho}-\frac{1}{\tau_{R}}{\mathbf{C}}(\Lambda)\vec{\rho}\,.$ (31)
@@ II.2.4 Computational Details
The matrix $\mathbf{F}$ gives free streaming dynamics and is dominant when $\tau/\tau_{R}\ll 1$. The second matrix $\mathbf{C}(\Lambda)$ gives the collision dynamics. The presence of $1/\tau_{R}\sim T^{1+\Lambda}$, makes the above differential equation nonlinear. However, the local (in time) dynamical behavior of the differential equation can be inferred from the eigenvalues of $\mathbf{C}(\Lambda)$. For late times, i.e $\tau/\tau_{R}>1$ we can assume the dynamics to be controlled by $\mathbf{C}(\Lambda)$. Then we can write a formal solution as,

### v5

@@ II.2.1 Basic definitions and properties
Multiplying Eq.(10) with $p^{n}P_{2l}({\text{cos}}\theta)$ and integrating over $dP$, we finally obtain the equation of motion governing the moment evolutions,
$\frac{\partial}{\partial\tau}\rho_{n,l}\penalty 10000\ +\penalty 10000\$$\frac{1}{\tau}\left[{\cal{P}}(n,l)\rho_{n,l-1}+{\cal{Q}}(n,l)\rho_{n,l}+{\cal{R}}(n,l)\rho_{n,l+1}\right]$
$=\int dP\left(p\right)^{n-1}P_{2l}(\text{cos}\theta){\cal{L}}[\phi_{p}]\penalty 10000\ ,$ (11)
with,
${\cal{P}}(n,l)=$ $2l\frac{(2l-1)}{(4l-1)}\frac{(n+2l)}{(4l+1)}\penalty 10000\ ,$
${\cal{Q}}(n,l)=$ $\frac{2}{3}+\frac{n(8l^{2}+4l-1)}{(4l-1)(4l+3)}+\frac{2l(2l+1)}{3(4l-1)(4l+3)}\penalty 10000\ ,$
${\cal{R}}(n,l)=$ $(n-2l-1)\frac{(2l+1)}{(4l+1)}\frac{(2l+2)}{(4l+3)}\penalty 10000\ .$ (12)
$(4l+1)xP_{2l}(x)=(2l+1)P_{2l+1}(x)+2l\penalty 10000\ P_{2l-1}(x)\penalty 10000\ ,$ (13)
$\partial P_{2l}(x)/\partial x=2l\left\{P_{2l-1}(x)-xP_{2l}(x)\right\}/(1-x^{2})\penalty 10000\ .$ (14)
$\frac{\partial}{\partial\tau}\rho_{n,l}\penalty 10000\ +\penalty 10000\$$\frac{1}{\tau}\left[{\cal{P}}\rho_{n,l-1}+{\cal{Q}}\rho_{n,l}+{\cal{R}}\rho_{n,l+1}\right]$
$=-\frac{1}{\tau_{R}}\left[\rho_{n,l}-\rho_{n,l}^{eq}\right]\penalty 10000\ ,$ (15)
II.2.2 Collision kernel with the momentum-dependent relaxation time approximation
The problem of a system with a momentum-dependent relaxation time has been investigated earlier using hydrodynamic equations [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. We investigate this from the perspective of kinetic theory, which gives insight into how such a system approaches equilibrium. Accordingly, while treating the collision term in the moment equation (11), we use the novel relaxation time collision operator [33] that conserves both particle number and energy-momentum irrespective of the momentum-dependence considered in $\tau_{R}(p)$ and the hydrodynamic frame chosen,
${\cal{L}}[\phi_{p}]=$ $-\frac{p}{\tau_{R}}f^{eq}\Bigg[\phi_{p}-P_{0}\frac{{\langle\frac{p}{\tau_{R}}P_{0}\phi_{p}\rangle}_{eq}}{{\langle\frac{p}{\tau_{R}}P_{0}P_{0}\rangle}_{eq}}$
$-P_{1}\frac{{\langle\frac{p}{\tau_{R}}P_{1}\phi_{p}\rangle}_{eq}}{{\langle\frac{p}{\tau_{R}}P_{1}P_{1}\rangle}_{eq}}-p^{\langle\mu\rangle}\frac{{\langle\frac{p}{\tau_{R}}p_{\langle\mu\rangle}\phi_{p}\rangle}_{eq}}{\frac{1}{3}{\langle\frac{p}{\tau_{R}}p_{\langle\nu\rangle}p^{\langle\nu\rangle}\rangle}_{eq}}\Bigg]\penalty 10000\ ,$ (16)
with $P_{0}=1$ and $P_{1}=1-p{\langle\frac{p}{\tau_{R}}\rangle}_{eq}/{\langle\frac{p^{2}}{\tau_{R}}\rangle}_{eq}$. The notation ${\langle\cdots\rangle}_{eq}$ reads, ${\langle\cdots\rangle}_{eq}=\int dPf_{p}^{eq}(\cdots)$. It should be mentioned here that a slightly different looking collision kernel is introduced in [41] based on the chosen momentum basis in which $\phi_{p}$ is expanded. However, it can be trivially seen that a momentum rearrangement proves their equivalence. Here, because of the neatness in the expression of collision term we proceed with ${\cal{L}}[\phi_{p}]$ given in (16).
With the given prescription, we denote the right-hand side of Eq.(11) (moment of the linearized collision term) as $C_{n,l,\Lambda}$ such that the moment evolution equation becomes,
$\penalty 10000\ \penalty 10000\ +\penalty 10000\ \penalty 10000\ C_{n,l,\Lambda}\penalty 10000\ .$
To proceed further, we decompose the momentum-dependent relaxation time $\tau_{R}(p)$ into a momentum-independent part $\tau_{R}$ (precisely a function of temperature $T$) and a part purely a function of single particle energy $p$ with the exponent $\Lambda$ being a positive number [30, 34],
$\tau_{R}(p)=\tau^{0}_{R}(T)p^{\Lambda}\penalty 10000\ .$ (18)
$\tau^{0}_{R}(T)=\left(\frac{\eta}{s}\right)\frac{5!}{\Gamma(5+\Lambda)}\frac{1}{T^{1+\Lambda}}\penalty 10000\ .$ (19)
From now on, whenever the notation $\tau_{R}$ has been mentioned either in the text or in the figures of the result section, it indicates this thermal part and must not be confused with the total relaxation time $\tau_{R}$.
Using expression (18), the moment of the collision term $C_{n,l,\Lambda}$ becomes (Appendix (A)),
$C_{n,l,\Lambda}=$ $-\frac{1}{\tau^{0}_{R}}\Bigg[\left\{\rho_{n-\Lambda,l}-\delta_{l0}\penalty 10000\ \rho^{eq}_{n-\Lambda,l}\right\}-A\penalty 10000\ \delta_{l0}\penalty 10000\ \rho^{eq}_{n-\Lambda,0}$
$-B\penalty 10000\ \delta_{l0}\penalty 10000\ \left\{\rho^{eq}_{n-\Lambda,0}-\frac{\rho^{eq}_{1-\Lambda,0}}{\rho^{eq}_{2-\Lambda,0}}\rho^{eq}_{n-\Lambda+1,0}\right\}\Bigg]\penalty 10000\ ,$ (20)
with,
$A=$ $\frac{\rho_{1-\Lambda,0}-\rho^{eq}_{1-\Lambda,0}}{\rho^{eq}_{1-\Lambda,0}}\penalty 10000\ ,$ (21)
$B=$ $\frac{\rho_{1-\Lambda,0}-\frac{\rho^{eq}_{1-\Lambda,0}}{\rho^{eq}_{2-\Lambda,0}}\rho_{2-\Lambda,0}}{\rho^{eq}_{3-\Lambda,0}\left(\frac{\rho^{eq}_{1-\Lambda,0}}{\rho^{eq}_{2-\Lambda,0}}\right)^{2}-\rho^{eq}_{1-\Lambda,0}}\penalty 10000\ .$ (22)
With the values of with $l=0$ and $n=1,2$ it can be observed that $C_{n,l,\Lambda}$ readily gives zero, preserving the collisional invariant property (conservation laws) independent of $\Lambda$. The resulting equations give the well-known conservation equations for particle number $({\rm n})$ and energy density $(\epsilon)$ for a system as follows,
$\frac{\partial{\rm n}}{\partial\tau}+\frac{{\rm n}}{\tau}=0\penalty 10000\ ,$ (23)
$\frac{\partial\epsilon}{\partial\tau}+\frac{4}{3}\frac{\epsilon}{\tau}=-\frac{2}{3}\frac{P_{L}-P_{T}}{\tau}\penalty 10000\ ,$ (24)
where we identify the moments as, $n=\rho_{1,0}\penalty 10000\ ,\penalty 10000\ \epsilon=\rho_{2,0}$ and $\rho_{2,1}=P_{L}-P_{T}=$ pressure anisotropy.
After some tedious algebra, the collision moment $C_{n,l,\Lambda}$ can be written in the following simplified form(Appendix (A)),
$C_{n,l,\Lambda}=$ $-\frac{1}{\tau_{R}^{0}}\Big[\rho_{n-\Lambda,l}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ +T^{n-2}K(n,2,\Lambda)C(n,\Lambda)\rho_{2-\Lambda,l}\Big\}\Big]\penalty 10000\ ,$ (25)
with the following two functions introduced,
$K(n,m,\Lambda)=\frac{\Gamma(n-\Lambda+2)}{\Gamma(m-\Lambda+2)}\penalty 10000\ ,$ (26)
$C(n,\Lambda)=\frac{1-K(1,n,\Lambda)K(n+1,2,\Lambda)}{1-K(1,2,\Lambda)K(3,2,\Lambda)}\penalty 10000\ ,$ (27)
such that $C(1,\Lambda)=0$, $C(2,\Lambda)=1$ and $K(n,n,\Lambda)=1$. This makes energy and number conservation explicit as $C_{n,l,\Lambda}$ is zero for $n=1,2$ and $l=0$. Here, the temperature $T$ and chemical potential $\mu$ is defined via the Landau conditons $\epsilon=\epsilon_{\text{eq}}$ and ${\rm n}={\rm n}_{\text{eq}}$,
${\rm n}$ $=e^{\mu/T}\frac{T^{3}}{2\pi^{2}}\Gamma(3)\,$, (28)
$\epsilon$ $=e^{\mu/T}\frac{T^{4}}{2\pi^{2}}\Gamma(4)\,.$ (29)
When $\Lambda$ is zero, the moments are coupled to each other only through the $l$ indices while for any non-zero $\Lambda$, each $(n,l)$ moment is coupled to the corresponding $(n-\Lambda,l)$ moment via the collision term $C_{n,l,\Lambda}$ (see Eq.(25)). This creates an infinite ladder of coupling towards the lower $n$ moments.
$\text{For\penalty 10000\ \penalty 10000\ }\penalty 10000\ l\neq 0\penalty 10000\ ,$
$\frac{\partial}{\partial\tau}\rho_{n,l}+\frac{1}{\tau}\left[{\cal{P}}\rho_{n,l-1}+{\cal{Q}}\rho_{n,l}+{\cal{R}}\rho_{n,l+1}\right]=-\frac{1}{\tau^{0}_{R}}\rho_{n-\Lambda,l}\penalty 10000\ ,$ (30)
$\text{For\penalty 10000\ \penalty 10000\ }\penalty 10000\ l=0\penalty 10000\ ,$
$\penalty 10000\ \penalty 10000\ -\frac{1}{\tau^{0}_{R}}\Big[\rho_{n-\Lambda,0}-T^{n-1}K(n,1,\Lambda)\left\{1-C(n,\Lambda)\right\}\rho_{1-\Lambda,0}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ -T^{n-2}K(n,2,\Lambda)C(n,\Lambda)\rho_{2-\Lambda,0}\Big]\penalty 10000\ .$ (31)
Eq.(30) is structurally not very different from the momentum-independent case (15) apart from the $\Lambda$ factor inclusion in the $n$ index of the moment to the source term at the right hand side (however, that change the dynamics anyway). But it is the $l=0$ moment in (31) that bears the effect of MDRTA the most via the $C(n,\Lambda)$ and $K(m,n,\Lambda)$ functions. It again carries its effect recursively through the couplings of the moment hierarchy as mentioned earlier.
III Computational Details
III.1 Initial Conditions
We consider a conformal system with particle mass $m=0$ at an initial temperature $T(\tau_{0})=1$ GeV and initial chemical potential $\mu(\tau_{0})=0$. We initialize the moments $\rho_{n,l}(\tau_{0})$ to their equilibrium values $\rho^{eq}_{n,l}$. Unlike momentum-independent relaxation time, systems with larger viscosity cannot be probed by considering smaller values of initial rescaled time. In this study, we choose two values of the initial viscosity to entropy ratio $\eta_{0}/s_{0}=0.2,0.02$. To make a comparison with the previous works [27], we then choose the initial time $\tau_{0}$ such that the scaled time is $\tau_{0}/\tau_{R}(\tau_{0})\in\{0.10,\penalty 10000\ 1.0\}$ so that the initial parameter set matches the case for $\Lambda=0$. We then solve the coupled equations in (17) by truncating the moment equations at $(n_{\rm min},n_{\rm max})=(0,5)$, and $(l_{\text{min}},l_{\text{max}})=(0,100)$ using the RK4 algorithm.
III.2 Truncation: Free Streaming
The moment evolution equation (17) is a set of infinitely coupled non-linear differential equations. To compute the moments, for practical purposes we need to truncate these equations at some finite order. In a previous work by de Brito et.al, [27] the $l$ moments were truncated by setting $\rho_{n,l}=0$ for some $l\geq l_{max}$, note that due to momentum independent relaxation time there was no $n$ coupling. This assumes that the contribution of the $l$ moments beyond $l_{max}$ can be ignored compared to the lower $l$ moments. But this type of truncation scheme no longer holds true for later times, when the $l$ moments decay and become comparable in value with respect to each other. The situation is elaborated below by analyzing different terms of Eq.(17). The free streaming evolution equation is given by,
$\frac{\partial}{\partial\tau}\rho_{n,l}=$ $-\frac{1}{\tau}{\cal{Q}}(n,l)\penalty 10000\ \rho_{n,l}$
$-\frac{1}{\tau}\left[{\cal{P}}(n,l)\penalty 10000\ \rho_{n,l-1}+{\cal{R}}(n,l)\penalty 10000\ \rho_{n,l+1}\right]\,.$ (32)
@@ III.2 Truncation: Free Streaming
$P(n,l\gg 1)$ $\sim\frac{l}{2}+\frac{(n+1)}{4}$ (33)
$Q(n,l\gg 1)$ $\sim\frac{(2n+3)}{4}$ (34)
$R(n,l\gg 1)$ $\sim-\frac{l}{2}+\frac{(n-2)}{4}\,.$ (35)
At late time we have $\rho_{n,l+1}\sim\rho_{n,l-1}$, and eventually the last two terms on the right-hand side of Eq.(32) partially cancel each other and for a large enough $\tau$. Therefore a better truncation for the free streaming evolution can be obtained by setting $\rho_{n,l_{\text{max}}-1}=\rho_{n,l_{\text{max}}+1}$.
Figure 1: The ratio of absolute values of alternate moments at late time for $n=2$ for a free streaming system. The ratio saturates near 1.
III.3 Truncation: MDRTA
Now, let us consider the case with a non-zero collision kernel. The truncation for the momentum-dependent is non-trivial as there is coupling in both $l$ and $n$ indices (see (25)). A non-zero $\Lambda$ fundamentally changes the type of the coupled differential equation. To understand this, let us rewrite the set of coupled differential equations in the following matrix representation
@@ III.3 Truncation: MDRTA
where $\vec{\rho}$ is a vector consisting of the infinite (or finite for a truncated system) set of $(n,l)$ moments and $\mathbf{F}$ and $\mathbf{C}(\Lambda)$ are matrices. The vector $\vec{\rho}$ is a collection of moments $\rho_{n,l}$, $n_{min}<n<n_{max}$, $0<l<l_{max}$, which are given some arbitrary ordering and written in a vector form $\vec{\rho}=(\rho_{0,0},\rho_{0,1}\dots,\rho_{1,0},\rho_{1,1}\dots)$. The matrices $\mathbf{F}$ and $\mathbf{C}(\Lambda)$ couples these moments. Specifically $\mathbf{F}$ couples the $l$ moments for a fixed $n$, while $\mathbf{C}(\Lambda)$ couples the $n$ moments(See Appendix (B.1) ).
@@
The matrix $\mathbf{F}$ gives free streaming dynamics and is dominant when $\tau/\tau_{R}\ll 1$. The second matrix $\mathbf{C}(\Lambda)$ gives the collision dynamics. The presence of $1/\tau_{R}\sim T^{1+\Lambda}$, makes the above differential equation nonlinear. However, the local (in time) dynamical behavior of the differential equation can be inferred from the eigenvalues of $\mathbf{C}(\Lambda)$. For late times, i.e $\tau/\tau_{R}>1$, we can assume the dynamics to be controlled by $\mathbf{C}(\Lambda)$. When the collision term is dominant, we can ignore the free streaming term $\mathbf{F}/\tau$ and write a formal solution to Eq.(36) as,
$\vec{\rho}_{n}(\tau)\sim\exp\left[{-\int d\tau\frac{1}{\tau^{0}_{R}}\mathbf{C}(\Lambda)}\right]\vec{\rho}_{n}(\tau_{0})\penalty 10000\ ,$ $\tau/\tau_{R}\gg 1.$ (37)
For $\Lambda=0$, the matrix is diagonal ($\mathbf{C}(\Lambda)\sim I$) and eigenvalues are positive and give rise to a pure decay. However, for non-zero $\Lambda$, $\mathbf{C}(\Lambda)$ (see Eq.(38) for n = 4 truncation) is off-diagonal and is similar to the left shift operator,
$\mathbf{C}(\Lambda)\sim\begin{bmatrix}0&&0&&0&&0\\ c_{n-4\Lambda}&&0&&0&&0\\ 0&&c_{n-3\Lambda}&&0&&0\\ 0&&0&&c_{n-2\Lambda}&&0\end{bmatrix}\,$. (38)
For finite truncations, the spectrum of the operator $\mathbf{C}(\Lambda)$, is no longer positive definite, and a late-time decay is not guaranteed. In Fig.(2), the evolution of temperature for two different values of $\Lambda\in\{0,0.125\}$ are given. We see that the temperature evolution diverges from the expected late-time decay behavior for $\Lambda=0.125$.
Figure 2: The evolution of temperature for $\Lambda=0$ and $\Lambda=0.125$. We see a growth in temperature when $\tau/\tau_{R}>1$ .
As the evolution of the system at late times is governed by the eigenvalues of the matrix $\mathbf{C}(\Lambda)$, we can understand the divergent behavior by analyzing its spectrum. As $\textbf{C}(\Lambda)$ is lower triangular, all its eigenvalues are zero and, therefore, $\mathbf{C}^{N}(\Lambda)=0$ (nilpotent), where $N$ is the dimension of the matrix. If $\mathbf{C}(\Lambda)$ is independent of time, the solution to Eq(37) is,
$\vec{\rho}(\tau)=\left(\sum_{k=0}^{N-1}\mathbf{C}(\Lambda)^{k}\tau^{k}\right)\vec{\rho}(\tau=\tau_{0})$ (39)
and therefore has polynomial growth instead of the expected exponential decay. This is the origin of the temperature growth seen in Fig.(2).
III.3.1 Moment Closure
We know that the linearized collision kernel L for the Boltzmann equation,
$p^{\mu}\partial_{\mu}f=-\textbf{L}(f-f^{eq})\,$, (40)
is positive semi-definite [46] and the zero eigenvalues of the linearized operator correspond to the collision invariants or the conserved quantities. This effectively enshrines the approach to equilibrium and entropy generation in the system. We expect this property to be preserved for any moment scheme representing the Boltzmann equation.
The divergence of temperature in Fig.(2) is due to improper truncation and closure of the moments. A proper truncation scheme should capture the salient features of the infinite hierarchy of moments. To recover the positive definiteness of the collision kernel, we use the Grad moment closure procedure. Grad’s procedure involves approximating the distribution function using a series expansion in momentum,